1. A solenoid consists of 200 turns of wire and has a length of 0.15m. What is the magnetic field inside

the solenoid if the current through it is a.) 0.24A, b.) 0.50A and c.) 1.2A


2. A circular wire loop with an area, A = 0.1m2 , is placed in magnetic field of magnitude B = 0.25 T.

Calculate the magnetic flux through the loop when a.) the plane of the loop is perpendicular to the field,

and b.) the normal plane of the loop makes an angle of 300 with respect to the field.


1
Expert's answer
2022-04-11T12:03:05-0400

1.

"B=\frac{\mu_0NI}l,"

"B_a=0.4~mT,"

"B_b=0.8~mT,"

"B_c=2~mT."

2.

"\Phi=AB\sin\alpha,"

"\Phi_a=25~mWb,"

"\Phi_b=13~mWb."